# Establishing a Connection Between the Jet and the Corona in Black Hole Low Mass X-Ray Binaries

### v1

@@ 1 Introduction
Black Hole Low-Mass X-ray Binaries (LMXBs) play an important role in understanding the universal phenomenon of accretion/ejection. During outburst cycles, LMXBs evolve rapidly compared to their more massive analogs (e.g. supermassive black holes) (e.g., Tetarenko et al., 2016; Corbel et al., 2013), making them ideal laboratories to study the accretion/ejection phenomena.
@@
As their outburst cycle progresses, LMXBs show characteristic X-ray spectral states, namely hard (HS), soft (SS), and intermediate states (IS) (e.g., Remillard & McClintock, 2006; Belloni, 2010). LMXB outbursts typically begin in the HS and transition into the SS via the IS before evolving back into the HS through a second IS. These states are thought to be connected to the geometry of the accretion flow near the black hole (e.g., Meyer-Hofmeister et al., 2009). During a typical outburst, LMXBs launch two types of jets: compact jets and transient jets, which are distinguished based on their morphology and radio spectrum. During the HS, partially self-absorbed compact jets produce a flat-to-inverted radio to mm spectrum (Hjellming & Johnston, 1988; Harmon et al., 1995; Fender, 2001, 2006; Gallo et al., 2004; Russell et al., 2013, 2014; Tetarenko et al., 2015), such that the spectral index, , is positive, if , where is flux density and is frequency. Near the X-ray peak of the outburst, as the source transitions from HS to the SS via the IS, the compact jet emission is observed to quench (Fender et al., 2004a; Russell et al., 2020), giving way to bright radio flaring, corresponding to the launching of the transient jet (Hjellming & Rupen, 1995; Fender et al., 1999; Fender, 2001; Carotenuto et al., 2021). The transient jets are composed of discrete knots of synchrotron emitting plasma that are launched out along the jet axis at close to the speed of light (Mirabel & Rodríguez, 1994). The transient jets exhibit an optically thin radio spectrum, where (Fender, 2006). Radio flaring is thought to arise as the ejecta interacts with the pre-existing compact jet or the surroundings (Fender, 2006).
@@
In the X-ray band, during the HS and IS, we generally observe back-scattered X-ray emission from the accretion disk, the so-called relativistic X-ray reflection (e.g., Lightman & White, 1988; Ross & Fabian, 1993; Bharali et al., 2019). The illuminating source creating the X-ray reflection is commonly referred to as the “corona”, which may be located in the inner part of the disk close to the black hole. The corona is believed to be formed by magnetically heated electrons, with a mean energy of 50 – 100 keV (e.g., Haardt & Maraschi, 1991, 1993). The accretion disk photons gain energy from the corona via inverse Compton scattering (Haardt & Maraschi, 1991, 1993). Despite significant efforts, the location and true nature of the corona remains ambiguous. Constraints have been placed on the coronal size, mainly from X-ray timing observations, finding the corona to be very compact and located close to the black hole (Kara et al., 2016).
@@
Early work linking the corona and jet by Henri & Pelletier (1991) suggested a two-phase model in which material ejected from the accretion disk heats the corona and accelerates it to relativistic speeds, producing the jets observed from AGN. Subsequent work on dynamic LMXB systems has continued to link the corona and jet. Vadawale et al. (2003) proposed that the ejection of coronal material may give rise to the transient jets. Markoff et al. (2005) found that the jets could produce some of the emissions attributed to the corona. Koljonen et al. (2015) observed a correlation between the X-ray photon index and properties of the radio jet, suggesting a causal connection between the corona and jet. From X-ray reflection fraction analysis, You et al. (2021) postulated that the corona is a standing shock with material flowing through it. Through analysis of the jet and X-ray spectra of three LMXBs, Cao et al. (2022) linked the softening of the X-ray spectrum to changes in the corona-jet system. Zhang et al. (2023) analyzed the energy and power density spectra of the BH-LMXB MAXI J1535–571, finding a vertically extended jet-like corona that contracts before the disappearance of the jet and begins to expand in its absence.
@@
The process by which particles are ejected via jets in LMXBs is not fully understood, complicating our study of the jet-corona relationship. In the case of compact jets, particles are accelerated into a power-law energy distribution through mechanisms such as shocks or magnetic reconnection (Markoff et al., 2001a; E. M. de Gouveia Dal Pino & Lazarian, A., 2005; de Gouveia Dal Pino, E. M. et al., 2010). This typically occurs in the jet launching region or at internal shock fronts along the jet (Tetarenko et al., 2019; Fender et al., 2004b; Malzac, 2014, 2012). Vadawale et al. (2003) proposed that an ejection of the corona during the transition to the SS could interact and shock the compact jet and lead to the creation of transient jets. However, the onset of compact jet quenching has been observed during hard-only outbursts, which fail to enter the soft state and a transient jet was not launched (e.g., Russell et al., 2014, 2015). Nevertheless, it remains possible that the compact jet is quenched by the transient jets, which are created by the ejected corona (Russell et al., 2020).
@@ 2 Observations and Methods
For our targets in this study, we require both X-ray spectral data and radio observations to track the evolving corona and jet. We focus our efforts on the detailed spectral modeling of the recent X-ray data and rely on published radio data to study the corresponding jet features. As NuSTAR’s 3.0-78.0 keV spectral range provides good coverage of the reprocessed spectral features, archival NuSTAR observations form the basis of our X-ray spectral analysis, with each NuSTAR observation forming an “epoch”. Where possible, we supplement the NuSTAR data with NICER or Swift/XRT data to provide coverage of the soft X-ray energy range and better constrain the disk and reflection features. For epochs with soft X-ray data, we only use the parts of the observations that occur simultaneously. Epoch dates, spectral states, and observations utilized are given in Table LABEL:results_table.
@@ 2.1 X-ray Data Reduction
All NICER data were reduced using NICERDAS version 2023-08-22_V011a and nicaldbver_xti_20221001. Cleaned event files were created from NICER data using . Source spectra and responses were created from the clean event files using the command with the specified overlapping GTIs. The background was estimated using the 111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ background model described in Remillard et al. (2022). We use the NICER spectral data in the keV range to avoid several well-studied instrumental and interstellar medium features below this range222https://heasarc.gsfc.nasa.gov/docs/nicer/data_analysis/workshops/NICER-CalStatus-Markwardt-2021.pdf.
For the Swift/XRT data analysis, we used xrtpipeline version 0.13.7 to generate clean event files. Further, we employed xselect to trim the event files to times concurrent with NuSTAR observations. Pileup was expected for many of the Swift/XRT spectra due to the brightness of these sources. To account for this, we used the prescription given by Romano et al. (2006) to remove the brightest source pixels.
@@ 2.1.1 MAXI J1535–571 Excluded X-ray
Data Initial modeling of our MJD 58010.26 observations of J1535 led to large residuals in the NuSTAR and particularly in NICER data. This epoch possesses three overlapping NuSTAR and NICER orbits. After inspecting the light curve of this observation, we found a sudden increase in the count rates for the two telescopes during the second orbit. While this flare may indeed be real, the study of such flares is not the focus of this paper. Thus, for this epoch, we only used data from the first and third overlapping orbits; this eliminated the residuals. Additionally, large residuals were present for the NICER data from MJD 58016.87 to MJD 58034.26 and were not used. Photon pileup could be the reason behind this because pile-up can become an issue in NICER above Crab (Center, 09/25/2024). XSPEC calculations of the NICER 2–10 keV flux reveal values significantly above this threshold for these epochs. Various studies reported that J1535 was extraordinarily bright in the soft X-ray band during these epochs (e.g., Tao et al., 2018).
@@ 2.2 X-ray Spectral Modeling
The geometry of the corona–disk is often inferred from the “reflection” spectrum (Reynolds & Nowak, 2003). Assuming a the corona is a compact source located at some height above the black hole (i.e., a lamp-post geometry), changes to this component of the X-ray spectrum can be connected to the corona’s distance (or height) from the black hole (Dauser et al., 2013). To constrain this height, we model the spectra of each outburst with three components. Firstly, a multi-colour disk blackbody component, described by the diskbb model (Mitsuda et al., 1984; Makishima et al., 1986). Secondly, we account for the primary X-ray continuum and relativistic reflection from the disk by including the lamp-post relxilllp model version 2.2 (Dauser & Garcia, 2022). Finally, galactic absorption along the line-of-sight is accounted for by the tbabs model (Wilms et al., 2000a). To account for any uncertainty in the cross-instrument calibration, we also include a constant model in our modeling that we freeze at one for NuSTAR/FPMA. All modeling was performed using XSPEC version 12.13 (Arnaud, 1996) with wilm abundances (Wilms et al., 2000b) and xsect set to vern (Verner et al., 1996).
@@
Multiple epochs are modeled together within XSPEC to constrain the model parameters better. For J1348, we model all epochs together. For J1820, we model our epochs in 3 separate groups, with MJDs 58191.94-58225.30 (initial HS) constituting group 1, MJDs 58242.27-58327.05 (covering the end of the HS, the IS, and the beginning of the SS) comprising group 2, and MJDs 58343.61-58404.95 (final HS) making group 3. For J1535, we split our epochs into two groups. Here, group 1 covers the initial HS and most of the first IS and consists of MJDs 58004.35-58017.21. The remaining observations (MJDS 58023.11-58090.77) cover the transition back to the HS and then to the SS and form group 2.
@@
We consider the black hole spin, the iron abundance, the inclination angle of the disk, and the column density to be “global” parameters, which we leave free but tie together between all epochs modeled together.
@@
In the relxilllp model, we let the model return the primary continuum and the expected reflection spectrum for a lamp post geometry. The disk temperature, disk normalization, corona height (), inner-disc radius, the power law index, the disk ionization, the cutoff energy, and relxilllp normalization were all left free to vary between epochs but were linked between observations within a given epoch.
@@
We estimated the errors of the best-fit parameters using the Markov Chain Monte Carlo (MCMC) simulations. For each group of epochs, we use XSPEC’s built-in Goodman-Weare MCMC algorithm with 270 walkers and a chain length of 6 million elements with the first 2 million elements burned. Errors for each parameter were estimated from the resulting chains by sorting the chain values and taking the central 90%. The numerical results of our X-ray spectral modeling can be found in Table LABEL:obstable.
@@ 2.2.1 MAXI J1820+070 Additional Blackbody Component
Our initial model failed to reproduce the observed spectra in several J1820 epochs. Excess residuals between our data and this model were observed in both NuSTAR and NICER spectra below keV our observations on MJDs 58201.53, 58201.86, 58224.96, 58259.33, 58297.20, 58315.02, 58349.82, and 58389.35. Following the analysis of the same data by Fabian et al. (2020), we add a blackbody component to our model for this source. This extra blackbody component likely arises from disk material emitting from or within the innermost stable circular orbit (ISCO), e.g. the plunging region (Mummery et al., 2024; Fabian et al., 2020). While a more robust model accounting for material within the plunging region has been developed by Mummery et al. (2024), a complete analysis with such a model is beyond the scope of this paper.
@@ 2.3 Analysis of Jet Features
We use the radio fluxes from prior work for all three sources to track the jet behavior. For J1348 and J1535, we also use the radio flux densities and spectral indices, , presented by Carotenuto et al. (2021) and Russell et al. (2019), respectively. For J1820, radio spectral indices were calculated by fitting a power law to the quasi-simultaneous radio flux values (taken from a variety of radio telescopes in the 0.1365-350 GHz range) reported by (Broderick et al., 2018; Bright et al., 2020; Echiburú-Trujillo et al., 2024; Shaw et al., 2021; Atri et al., 2020)and compiled in Echiburú-Trujillo et al. (2024). Throughout this work, we interpret a shift from optically thick to thin values of as a quenching of the compact jet, as it would be expected as part of a transition to the transient jet. We additionally use prior work to determine the dates of the launching of discrete jet ejecta. Jet launch dates for J1348, J1535, and J1820 are taken from Carotenuto et al. (2021), Russell et al. (2019), and Wood et al. (2021) respectively.

### v2

@@ 1 Introduction
Black Hole Low-Mass X-ray Binaries (LMXBs) play an important role in understanding the universal phenomenon of accretion/ejection. During outburst cycles, LMXBs evolve rapidly compared to their more massive analogues, supermassive black holes (e.g., 2004cbhg.symp..169M; hopkins2005), making LMXBs ideal laboratories to study the accretion/ejection phenomena.
@@
As their outburst cycles progresses, LMXBs show characteristic X-ray spectral states, namely the low-hard state (LHS), high-soft state (HSS), and the Intermediate State (IMS). The IMS can be further subdivided into the hard (HIMS) and soft (SIMS) intermediate states based primarily on spectral timing properties (e.g., Remillard2006; Belloni2010). LMXB outbursts typically begin in the LHS and transition through the HIMS and SIMS to the HSS before evolving back to the LHS via the IMS. These states are thought to be connected to the geometry of the accretion flow near the black hole (e.g., Meyer-Hofmeister2009). During a typical outburst, LMXBs launch two types of jets: compact jets and transient jets, which are distinguished based on their morphology and radio spectrum. During the LHS, partially self-absorbed compact jets produce a flat-to-inverted radio to mm spectrum (Hjellming1988; Harmon1995; Fender2001; Fender2006; gallo_2004; Russell_D_2013; Russell_2014_1836; Tetarenk_2016), such that the spectral index, , is positive, if , where is flux density and is frequency. Near the X-ray peak of the outburst, as the source transitions from HIMS to the SIMS, the compact jet emission is observed to quench (fend04; Russell_2020), giving way to bright radio flaring, corresponding to the launching of the transient jet (Hjellming95; Fender_1999; Fender2001; car). The transient jets are composed of discrete knots of synchrotron emitting plasma that are launched out along the jet axis at close to the speed of light (mirabel_2004). The transient jets exhibit an optically thin radio spectrum, where (Fender2006). Radio flaring is thought to arise as the ejecta interacts with the pre-existing compact jet or the surroundings (Fender2006).
@@
In the X-ray band, we generally observe back-scattered X-ray emission from the accretion disk, the so-called relativistic X-ray reflection (e.g., Lightman1988; Ross1993; Bharali2019). The illuminating source creating the X-ray reflection is commonly referred to as the “corona”, which may be located in the inner part of the disk close to the black hole. The corona is believed to be formed by magnetically heated electrons, with a mean energy of 50 – 100 keV (e.g., haardt91; Haardt1993). The accretion disk photons gain energy from the corona via inverse Compton scattering (haardt91; Haardt1993). Despite significant efforts, the location and true nature of the corona remains ambiguous. Constraints have been placed on the coronal size, mainly from X-ray reverberation studies, finding the corona to be compact and located close to the black hole (Kara2016). Other studies aimed at constraining the geometry of the corona have modeled the RMS-variability and phase-lags of quasi-periodic oscillations (QPOs) observed in LMXB spectra in addition to the time averaged spectra. Using spectral-timing models, Garcia2021 and bellavita2022 found that the corona in MAXI J1348–630 can be best described as consisting of two components one small and one more extended. A study of MAXI J1348–630’s QPOs during the rise of the outburst and reflares also suggests a two Component corona, with one being more horizontally extended (along the disk) corona to explain the high fraction emission reflected by the disk (alabarta2025). According to rawat2023, a two-coronal model is also preferred in the case of MAXI J1535–571. Studying the QPOs during the SIMS, zhang2023 suggests a vertically extended jet-like corona in the SIMS. This more complex and extended coronal geometry, is in agreement with recent results from (IXPE) ixpe, which have measured the X-ray polarization of several galactic black holes and thus provided evidence of a corona that is elongated and oriented along the disk (Veledina_2023_ixpe; cyg_ixpe, e.g.).
A changing coronal geometry has also been suggested, ma2023 studied QPOs in MAXI J1820+020 in its HIMS-SIMS transition, suggesting that the corona expands horizontally along the disk during the outburst rise and then transforms into a jet-like extended shape in the SIMS,accompanied by a radio flare. belloni2024 studied QPOs in GRS 1915+105 to find a rapidly varying corona—with its size varying between and in a single 10.5 hour observation for a BH reid2023.
@@
Early work linking the corona and jet by gilles1991 suggested a two-phase model in which material ejected from the accretion disk heats the corona and accelerates it to relativistic speeds, producing the jets observed from AGN. Subsequent work on dynamic LMXB systems has continued to link the corona and jet. Vadawale2003 proposed that the ejection of coronal material may give rise to the transient jets. markoff_2005 found that the jets could produce some of the emissions attributed to the corona. koljonen_2015 observed a correlation between the X-ray photon index and properties of the radio jet, suggesting a causal connection between the corona and jet. From X-ray reflection fraction analysis, you2021 postulated that the corona is a standing shock with material flowing through it. Through analysis of the jet and X-ray spectra of three LMXBs, cao2022 linked the softening of the X-ray spectrum to changes in the corona-jet system. A study of radio and X-ray data of GRS 1915+105 covering a 10 year period also revealed a correlation between the radio flux from the jet and the strength of the Fe K line Mendez2022, which is thought to arise from the X-ray reflection of coronal emission. Several studies performing specto-temporal analysis in the HIMS and SIMS of MAXI J1535–571 using Insight–HXMT and NICER observations have found a vertically extended jet-like corona that contracts before the disappearance of the jet and begins to expand in its absence (Zhang2022; zhang2023; rawat2023)
@@
The process by which particles are ejected via jets in LMXBs is not fully understood, complicating our study of the jet-corona relationship. In the case of compact jets, particles are accelerated into a power-law energy distribution through mechanisms such as shocks or magnetic reconnection (markoff_2001; dalpino_2005; dalpino_2010). This typically occurs in the jet launching region or at internal shock fronts along the jet (Tetarenko_2019; Fender2004; Malzac2014; Malzac_2012). Vadawale2003 proposed that an ejection of the corona during the transition to the HSS could interact and shock the compact jet and lead to the creation of transient jets. However, the onset of compact jet quenching has been observed during hard-only outbursts, which fail to enter the soft state and a transient jet was not launched (e.g., Russell_2014; Russell2015). Nevertheless, it remains possible that the compact jet is quenched by the transient jets, which are created by the ejected corona (Russell_2020).
@@ 2 Observations and Methods
We rely on prior work to determine the spectral states and the dates of transition between them. For J1820, we adopt the state transitions dates determined by Shidatsu_2019, and identify the HIMS–SIMS transition by the QPO change reported in (Homan2020). For J1348, we use the state transitions outlined in zhang, while for J1535 we take spectral states from Tao_2018 and nakahira_2018 compiled by Russell_2019.
For our targets in this study, we require both X-ray spectral data and radio observations to track the evolving corona and jet. We focus our efforts on the detailed spectral modeling of the recent X-ray data and rely on published radio data to study the corresponding jet features. As NuSTAR’s 3.0-78.0 keV spectral range provides good coverage of the reprocessed spectral features, archival NuSTAR observations form the basis of our X-ray spectral analysis, with each NuSTAR observation forming an “epoch”. Where possible, we supplement the NuSTAR data with NICER or Swift/XRT data to provide coverage of the soft X-ray energy range and better constrain the disk and reflection features. For epochs with soft X-ray data, we only use the parts of the observations that occur simultaneously. Epoch dates, spectral states, and observations utilized are given in Table 1.
@@ 2.1 X-ray Data Reduction
All NICER data were reduced using NICERDAS version 2023-08-22_V011a and nicaldbver_xti_20221001. Cleaned event files were created from NICER data using . Source spectra and responses were created from the clean event files using the command with the specified overlapping GTIs. The background was estimated using the 111https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/nicerl3-spect/ background model described in Remillard2022. We use the NICER spectral data in the keV range to avoid several well-studied instrumental and interstellar medium features below this range222https://heasarc.gsfc.nasa.gov/docs/nicer/data_analysis/workshops/NICER-CalStatus-Markwardt-2021.pdf.
For the Swift/XRT data analysis, we used xrtpipeline version 0.13.7 to generate clean event files. Further, we employed xselect to trim the event files to times concurrent with NuSTAR observations. Pileup was expected for many of the Swift/XRT spectra due to the brightness of these sources. To account for this, we used the prescription given by Romano2006 to remove the brightest source pixels.
@@ 2.1.1 MAXI J1535–571 Excluded X-ray
Data Initial modeling of our MJD 58010.26 observations of J1535 led to large residuals in the NuSTAR and particularly in NICER data. This epoch possesses three overlapping NuSTAR and NICER orbits. After inspecting the light curve of this observation, we found a sudden increase in the count rates for the two telescopes during the second orbit. While this flare may indeed be real, the study of such flares is not the focus of this paper. Thus, for this epoch, we only used data from the first and third overlapping orbits; this eliminated the residuals. Additionally, large residuals were present for the NICER data from MJD 58016.87 to MJD 58034.26 and were not used. Photon pileup could be the reason behind this because pile-up can become an issue in NICER above Crab (nicer_pileup). XSPEC calculations of the NICER 2–10 keV flux reveal values significantly above this threshold for these epochs. Various studies reported that J1535 was extraordinarily bright in the soft X-ray band during these epochs (e.g., Tao_2018).
@@ 2.2 X-ray Spectral Modeling
The geometry of the corona–disk is often inferred from the “reflection” spectrum (Reynolds2003). Approximating the corona as a compact source located at some height above the black hole (i.e., a lamp-post geometry), changes to this component of the X-ray spectrum can be connected to the corona’s distance (or height) from the black hole (dauser). While the corona is likely more extended, as indicated by recent polarization studies (see discussion), we adopt this simple lamp-post geometry as a probe of the evolution of the Comptonizing region for its ease of use and interpretation. To constrain this height, we model the spectra of each outburst with three components. Firstly, a multi-colour disk blackbody component, described by the diskbb model (diskbb; Makishima1986). Secondly, we account for the primary X-ray continuum and relativistic reflection from the disk by including the lamp-post relxilllp model version 2.2 (relxill). Finally, galactic absorption along the line-of-sight is accounted for by the tbabs model (tbabs). To account for any uncertainty in the cross-instrument calibration, we also include a constant model in our modeling that we freeze at one for NuSTAR/FPMA. All modeling was performed using XSPEC version 12.13 (xspec) with wilm abundances (wilm) and xsect set to vern (verners).
@@
Multiple epochs are modeled together within XSPEC to constrain the model parameters better. For J1348, we model all epochs together. Due to computational limitations arising from the large number of epochs in the other two sources, we modeled these epochs in separate groups rather than all together. For J1820, we model our epochs in 3 separate groups, with MJDs 58191.94-58225.30 (initial LHS) constituting group 1, MJDs 58242.27-58327.05 (covering the end of the LHS, the HIMS and SIMS, and the beginning of the HSS) comprising group 2, and MJDs 58343.61-58404.95 (final LHS) making group 3. For J1535, we split our epochs into two groups. Here, group 1 covers the initial LHS and most of the transition to the HSS and consists of MJDs 58004.35-58017.21. The remaining observations (MJDS 58023.11-58090.77) cover the transition back to the LHS and then to the HSS and form group 2.
@@
We consider the black hole spin, the iron abundance, the inclination angle of the disk, and the column density to be “global” parameters, which we leave free but tie together between all epochs in each group. These global parameters were allowed to vary between groups. We discuss the impact of this further in section 3.
@@
In the relxilllp model, we let the model return the primary continuum and the expected reflection spectrum for a lamp post geometry, while diskbb accounts for the temperature and normalization of the accretion disk. The disk temperature, disk normalization, corona height (), inner-disc radius, the power law index, the disk ionization, the cutoff energy, and relxilllp normalization were all left free to vary between epochs but were linked between observations within a given epoch.
@@
We estimated the errors of the best-fit parameters using the Markov Chain Monte Carlo (MCMC) simulations. For each group of epochs, we use XSPEC’s built-in Goodman-Weare MCMC algorithm with 270 walkers and a chain length of 6 million elements with the first 2 million elements burned. Errors for each parameter were estimated from the resulting chains by sorting the chain values and taking the central 90%. The numerical results of our X-ray spectral modeling can be found in Table 1.
@@ 2.2.1 MAXI J1820+070 Additional Blackbody Component
Our initial model failed to reproduce the observed spectra in several J1820 epochs. Excess residuals between our data and this model were observed in both NuSTAR and NICER spectra below keV our observations on MJDs 58201.53, 58201.86, 58224.96, 58259.33, 58297.20, 58315.02, 58349.82, and 58389.35. Following the analysis of the same data by Fabian_2020, we add a blackbody component to our model for this source. This extra blackbody component likely arises from disk material emitting from or within the innermost stable circular orbit (ISCO), e.g. the plunging region (1820_isco; Fabian_2020). While a more robust model accounting for material within the plunging region has been developed by 1820_isco, a complete analysis with such a model is beyond the scope of this paper.
@@ 2.3 Analysis of Jet Features
We use the radio fluxes from prior work for all three sources to track the jet behavior. For J1348 and J1535, we also use the radio flux densities and spectral indices, , presented by car and Russell_2019, respectively. For J1820, radio spectral indices were calculated by fitting a power law to the quasi-simultaneous radio flux values (taken from a variety of radio telescopes in the 0.1365-350 GHz range) reported by (atel11887; Bright_2020; Echibur_Trujillo_2024; Shaw_2021; atri2020) and compiled in Echibur_Trujillo_2024. Throughout this work, we interpret a shift from optically thick to thin values of as a quenching of the compact jet, as it would be expected as part of a transition to the transient jet. We note that although the compact jet is quenched during the HSS, steep-spectrum radio emission can linger at the core of the outburst deep into the HSS (e.g., Russell_2019; hughes_2025a). For J1535, this core radio emission faded 5 months into the HSS, while in the case of the black hole low-mass X-ray binary Swift J1727.8–1613, the radio emission from the core remained until the transition back to the LHS (hughes2025b). This lingering emission is often attributed to transient ejecta that have not traveled far enough to be resolved from the core position.
To determine the dates of the launching of discrete jet ejecta we rely on values reported in the literature. Jet launch dates for J1535, and J1820 are taken from Russell_2019 and Wood_2021 respectively. In the case of MAXI J1348-630, there were multiple excursions to the IMS during the HSS (car; zhang). One of these excursions launched a second ejection event that became resolved from the compact source. car detected both the first and second ejection events as they propagated away from the compact object.